Its 6 over 5x to the power of 8 plus 4 over 3x to the power of 6
Answer by edjones(8007)

6/5x^8 + 4/3x^6 (5x^8*3x^6=15x^14=LCD)
=(18x^6+20x^8)/15x^14
=x^6(20x^2+18)/15x^14
=(20x^2+18)/15x^8
.
Ed